### v1

# High-dimensional long-range statistical mechanical models have random walk correlation functions

Abstract
We consider long-range percolation, Ising model, and self-avoiding walk on $\mathbb{Z}^{d}$, with couplings decaying like $|x|^{-(d+\alpha)}$ where $0<\alpha\leq 2$, above the upper critical dimensions. In the spread-out setting where the lace expansion applies, we show that the two-point function for each of these models exactly coincides with a random walk two-point function, up to a constant prefactor. Using this, we prove a uniform upper bound of the form $|x|^{-(d-\alpha)}\min\{1,(p_{c}-p)^{-2}|x|^{-2\alpha}\}$ (with logarithmic corrections if $\alpha=2$) for the two-point function near the critical point $p_{c}$. We also give a simple proof of the convergence of the lace expansion, assuming diagrammatic estimates.
1 Introduction and results
1.1 Introduction
Statistical mechanical models on $\mathbb{Z}^{d}$ exhibit mean-field behaviours above an upper critical dimension. Models with short-range (finite-range or exponentially decaying) couplings have been studied extensively in the literature, and the upper critical dimension is $d_{c}=6$ for percolation and is $d_{c}=4$ for the Ising model and the self-avoiding walk (SAW). Two main tools to prove mean-field behaviours for percolation and the self-avoiding walk are the lace expansion [6, 17, 30, 16, 15] and the (reversed) Simon–Lieb inequality [11, 10]. The Ising model offers a wider range of tools, and mean-field results in dimensions $d>d_{c}$ have been obtained using the lace expansion [28], correlation inequalities [13], and methods based on the random current representation [12, 1, 3]; the last is also used to prove the triviality of the Ising model at $d=d_{c}=4$ [2].
Long-range models have couplings that decay polynomially like $\lvert x\rvert^{-(d+\alpha)}$ with $\alpha>0$. These models have upper critical dimension
$d_{c}=d_{c}(\alpha)=\begin{cases}2\times\min\{\alpha,2\}&(\text{Ising, SAW})\\ 3\times\min\{\alpha,2\}&(\text{percolation}),\end{cases}$ (1.1)
and they can exhibit mean-field behaviours in low dimensions if $\alpha>0$ is sufficiently small. Both the lace expansion [18, 8, 9] and the random current method for the Ising model [27] have been extended to this setting. For percolation, a different non-perturbative approach to mean-field behaviours is developed recently in [19].
Common in the analyses based on the lace expansion or the Simon–Lieb inequality is the idea that an interacting model in dimensions $d>d_{c}$ should be close to a random walk (a free model). Most previous works compare the model to the simple random walk on the underlying graph; [6, 17, 30, 18] used such comparisons in the Fourier space, and [16, 28, 8, 11, 10] used such comparisons in the physical space. Despite a simple idea, the freedom to choose a random walk to compare to has not been explored much. Two exceptions are [9, 22], where suitable random walks are chosen based on the interacting model to ease the analyses. In this paper, we push the comparison idea to the extreme, by showing that the two-point function $G_{p}$ for spread-out long-range models exactly coincides with the two-point function $S_{\mu_{p}}$ of some random walk, with a perturbed one-step distribution, when $\alpha\leq 2$:
$G_{p}(x)=A_{p}S_{\mu_{p}}(x)\qquad(x\in\mathbb{Z}^{d})$ (1.2)
for some constant $A_{p}$. The proof uses the lace expansion and is inspired by the random walk used in [9]. The main idea is to include the effects of interaction (the lace function), which decays faster than $\lvert x\rvert^{-(d+\alpha)}$ when $\alpha$ is small, into the one-step distribution.
Using the exact random walk representation, the proof of convergence of the lace expansion avoids deconvolution theorems as in [16, 25, 8, 9] and is arguably one of the simplest. Interestingly, when $\alpha$ is larger so that it is no longer possible to include all effects of interaction into the one-step distribution, the simple deconvolution method of [26, 25] can be applied. The new method we develop in this paper therefore complements [25] and covers all remaining values of $\alpha>0$.
We also use the random walk representation to prove a uniform upper bound for the subcritical two-point function $G_{p}$ near the critical point $p_{c}$, of the form
$G_{p}(x)\leq\frac{C}{\lvert x\rvert^{d-\alpha}}\min\biggl{\{}1,\frac{1}{(p_{c}-p)^{2}\lvert x\rvert^{2\alpha}}\biggr{\}},$ (1.3)
with logarithmic corrections if $\alpha=2$. The upper bound interpolates between the $\lvert x\rvert^{-(d-\alpha)}$ behaviour at criticality [8, 9] and the ${\chi(p)}^{2}\lvert x\rvert^{-(d+\alpha)}$ behaviour below criticality [4], where ${\chi(p)}=\sum_{x\in\mathbb{Z}^{d}}G_{p}(x)\asymp(p_{c}-p)^{-1}$ is the susceptibility. This is in contrast with short-range models, for which subcritical two-point functions decay exponentially and (1.3) should take the form
$G_{p}(x)\leq\frac{C}{|x|^{d-2}}\exp\Bigl{\{}-c(p_{c}-p)^{1/2}|x|\Bigr{\}}.$ (1.4)
As discussed in [24], the uniform bound (1.3) near criticality can be used to study long-range models on the $d$-dimensional discrete torus when $d>d_{c}$, and it can be used to prove that the torus two-point function levels off to a nonzero constant called the “torus plateau.” For short-range models, both (1.4) and the torus plateau have been proved for percolation [20], the Ising model [23], and the self-avoiding walk [21, 31]. For long-range models, (1.3) is only known for percolation when $\alpha<1$ and $d\geq 1$ [19], with a matching lower bound; the lower bound is also known for the Ising model under the same conditions [27, Proposition 3.25]. Our method yields the upper bound (1.3) for all $\alpha\leq 2$.
Notation. We write $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. We write $f\lesssim g$ to mean that there is a $C>0$ such that $f\leq Cg$, write $f\asymp g$ to mean that $f\lesssim g$ and $g\lesssim f$, and write $f\sim g$ to mean that $\lim f/g=1$. With $L\geq 1$ and $|x|$ the Euclidean norm of $x\in\mathbb{R}^{d}$, we define
$|\!|\!|x|\!|\!|_{L}=\frac{\pi}{2}\max\{\lvert x\rvert,L\}.$ (1.5)
Note that (1.5) does not define a norm on $\mathbb{R}^{d}$ but satisfies $|\!|\!|x|\!|\!|_{L}=L|\!|\!|\frac{x}{L}|\!|\!|_{1}$. Also note that $\log|\!|\!|\frac{x}{L}|\!|\!|_{1}\geq\log\frac{\pi}{2}>0$ for all $x$.
1.2 The models
We consider percolation, the Ising model, and the self-avoiding walk with long-range coupling, described by a probability distribution $D$ on $\mathbb{Z}^{d}$ that satisfies the following assumption. The spread-out parameter $L\geq 1$ will be taken large.
Assumption 1.1.
Let $L\geq 1$ and $\alpha>0$. We assume $D:\mathbb{Z}^{d}\to[0,\infty)$ is a probability distribution that is $\mathbb{Z}^{d}$-symmetric (invariant under reflection in coordinate hyperplanes and rotation by $\pi/2$) and satisfies $D(0)=0$ and
$\frac{K_{D}^{-1}L^{\alpha}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}\leq D(x)\leq\frac{K_{D}L^{\alpha}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}\qquad(x\neq 0)$ (1.6)
for some constants $K_{D}\geq 1$. We also assume there is a constant $\Delta\in(0,1)$ for which the Fourier transform $\hat{D}(k)=\sum_{x\in\mathbb{Z}^{d}}D(x)e^{ik\cdot x}$ obeys
$1-\hat{D}(k)\begin{cases}<2-\Delta&(k\in(-\pi,\pi]^{d})\\ >\Delta&(\lVert k\rVert_{\infty}>L^{-1}).\end{cases}$ (1.7)
[…]
$D(x)=\frac{v(x/L)}{\sum_{x\neq 0}v(x/L)}\mathds{1}_{x\neq 0},$ (1.8)
then the denominator is asymptotic to $L^{d}$ by a Riemann-sum approximation, and (1.6) holds. Property (1.7) is verified in [7, Proposition 1.1] for $L$ sufficiently large.
We now define the models. Let $D$ obey Assumption 1.1. We write $\omega=(\omega_{0},\dots,\omega_{\lvert\omega\rvert})\in(\mathbb{Z}^{d})^{\mathbb{N}}$ for a path in $\mathbb{Z}^{d}$, where $\lvert\omega\rvert$ denotes the length of the path. For $x\in\mathbb{Z}^{d}$, we write $\mathcal{W}(x)$ for the set of all paths from 00 to $x$, and we write $\mathcal{S}(x)$ for the set of all self-avoiding paths from 00 to $x$, i.e., the set of paths $\omega$ such that $\omega_{i}\neq\omega_{j}$ for all $0\leq i<j\leq\lvert\omega\rvert$. Given $p\geq 0$ and $x\in\mathbb{Z}^{d}$, the random walk two-point function and the self-avoiding walk two-point function are defined respectively by
$C_{p}(x)=\sum_{\omega\in\mathcal{W}(x)}\prod_{j=1}^{\lvert\omega\rvert}pD(\omega_{j}-\omega_{j-1}),\qquad G_{p}^{\mathrm{(SAW})}(x)=\sum_{\omega\in\mathcal{S}(x)}\prod_{j=1}^{\lvert\omega\rvert}pD(\omega_{j}-\omega_{j-1}).$ (1.9)
The sum defining $C_{p}(x)$ converges for all $p\in[0,1)$, and it converges also at the critical point $p=1$ in dimensions $d>\alpha\wedge 2$. We will also consider a random walk with a different one-step distribution $P(x)$, whose two-point function we denote by $S_{p}(x)$.
For percolation, let $p\in[0,\lVert D\rVert_{\infty}^{-1}]$. To each pair of vertices $\{x,y\}$ in $\mathbb{Z}^{d}$, we associate independently a Bernoulli random variables $n_{\{x,y\}}$ with
$\mathbb{P}_{p}(n_{\{x,y\}}=1)=pD(x-y).$ (1.10)
A configuration is a realisation of these random variables. We say the bond $\{x,y\}$ is occupied if $n_{\{x,y\}}=1$ and is vacant otherwise. The set of all vertices $x$ that are connected to 00 by some path of occupied bonds is denoted $\mathcal{C}(0)$. The percolation two-point function is defined as
$G_{p}^{(\mathrm{perc})}(x)=\mathbb{P}_{p}(x\in\mathcal{C}(0)).$ (1.11)
For the Ising model, we assume a ferromagnetic coupling $J:\mathbb{Z}^{d}\to[0,\infty)$ that obeys Assumption 1.1 (replace $D$ by $J$), and we first consider a finite subset $\Lambda\subset\mathbb{Z}^{d}$ containing 00. The Ising model on $\Lambda$ is a family of probability measures on spin configurations $\sigma\in\{\pm 1\}^{\Lambda}$, defined using the Hamiltonian
$H^{\Lambda}(\sigma)=-\frac{1}{2}\sum_{x,y\in\Lambda}J(y-x)\sigma_{x}\sigma_{y}.$ (1.12)
At inverse temperature $\beta\geq 0$, the Ising two-point function is defined by the infinite-volume limit
$\tilde{G}_{\beta}^{(\mathrm{Ising})}(x)=\lim_{\Lambda\uparrow\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}^{\Lambda}=\lim_{\Lambda\uparrow\mathbb{Z}^{d}}\frac{\sum_{\sigma\in\{\pm 1\}^{\Lambda}}\sigma_{0}\sigma_{x}e^{-\beta H^{\Lambda}(\sigma)}}{\sum_{\sigma\in\{\pm 1\}^{\Lambda}}e^{-\beta H^{\Lambda}(\sigma)}},$ (1.13)
which exists because $\langle\sigma_{0}\sigma_{x}\rangle_{\beta}^{\Lambda}$ is increasing in $\Lambda$ by Griffiths inequalities. For $\beta>0$, it is convenient to define a probability distribution $D$ by
$D(x)=p^{-1}\tanh(\beta J(x)),\qquad p=\sum_{y\in\mathbb{Z}^{d}}\tanh(\beta J(y)).$ (1.14)
Note that $p$ is strictly increasing and continuous in $\beta$, so we can write $G_{p}^{(\mathrm{Ising})}(x)=\tilde{G}_{\beta}^{(\mathrm{Ising})}(x)$. It follows from $\lVert J\rVert_{\infty}\lesssim L^{-d}$ and from the expansion $\tanh(z)=z+O(z^{3})$ that
$\frac{p}{\beta},\,\frac{D(x)}{J(x)}=1+O(L^{-2d})$ (1.15)
uniformly in $x\neq 0$ and in $0<\beta\leq 4$, when $L$ is large. Thus, $D$ satisfies Assumption 1.1 with $\beta$-independent constants that are slightly perturbed from those of $J$.
We omit the model labels on $G_{p}$ from now on. For each of the models, there is a critical point $p_{c}\geq 1$ such that $\chi(p)=\sum_{x\in\mathbb{Z}^{d}}G_{p}(x)$ is finite if and only if $p\in[0,p_{c})$. Under slightly stronger assumptions on $D$, Chen and Sakai [8, 9] proved that, in dimensions $d>d_{c}$ (for $\alpha\neq 2$) or $d\geq d_{c}$ (for $\alpha=2$), the critical two-point function $G_{p_{c}}(x)$ is finite and behaves as $\lvert x\rvert\to\infty$ as
$G_{p_{c}}(x)\sim\frac{\mathrm{const}}{\lvert x\rvert^{d-\alpha\wedge 2}}\times\begin{cases}1&(\alpha\neq 2)\\ (\log\lvert x\rvert)^{-1}&(\alpha=2).\end{cases}$ (1.16)
In particular, this implies finiteness of the bubble diagram $B(p_{c})=G_{p_{c}}*G_{p_{c}}(0)$ for Ising and SAW and finiteness of the triangle diagram $T(p_{c})=G_{p_{c}}^{*3}(0)$ for percolation, so that the critical exponents $\gamma,\beta,\delta$ take their mean-field values. Note the logarithmic correction for $\alpha=2$ makes the models mean-field also at the critical dimension $d_{c}$.
1.3 Main results
Our main result is the following theorem, which represents the two-point function for percolation, the Ising model, or the self-avoiding walk as a random walk two-point function. Let
=$\ell=\begin{cases}3&(\text{Ising, SAW})\\ 2&(\text{percolation}).\end{cases}$ (1.17)
Theorem 1.2.
Let $0<\alpha\leq 2+(\ell-1)(d-d_{c})$. Let $d>d_{c}$ if $\alpha\neq 2$, and let $d\geq d_{c}$ if $\alpha=2$. Let $D$ (for percolation or SAW) or $J$ (for Ising) obey Assumption 1.1, with $L\geq L_{0}$ where $L_{0}$ is sufficiently large. Then for all $p\in[\frac{1}{2},p_{c}]$, there are $A_{p}>0$, $\mu_{p}\in[0,1]$, and a probability distribution $P$ depending on $p$ such that
$G_{p}(x)=A_{p}S_{\mu_{p}}(x)\qquad(x\in\mathbb{Z}^{d}),$ (1.18)
where $S_{\mu}$ denotes the random walk two-point function generated by $P$. Moreover, we have ${\mu_{p_{c}}}=1$ and
$A_{p},\,\frac{{\mu_{p}}}{p}=1+O(L^{-d}),\qquad\frac{P(x)}{D(x)}=1+O(L^{-(\ell-1)d})$ (1.19)
uniformly in $p\in[\frac{1}{2},p_{c}]$ and in $x\neq 0$.
Theorem 1.2 gives a way to study $G_{p}$ by studying a random walk with a slightly perturbed kernel $P$ that still obeys Assumption 1.1. We use it to prove a near-critical estimate for $G_{p}$. (The critical behaviour can be obtained from that of $S_{1}$, by, e.g., [9, Theorem 1.3]).) The next proposition is a near-critical estimate for random walks obeying Assumption 1.1. The critical $\mu=1$ case of the proposition has been proved in [8, 9], and we extend their arguments to $\mu<1$.
Proposition 1.3.
Let $\alpha>0$, $d>\alpha\wedge 2$, and let $D$ obey Assumption 1.1. If $\alpha\neq 2$, there is a constant $K_{C}>0$ depending only on $K_{D}$ and $\Delta$ such that for all $\mu\in[0,1]$ and $x\in\mathbb{Z}^{d}$,
$C_{\mu}(x)\leq\delta_{0,x}+\frac{K_{C}}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\bigg{(}1\wedge\frac{1}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{2(\alpha\wedge 2)}(1-\mu)^{2}}\bigg{)}.$ (1.20)
If $\alpha=2$, we instead have
$C_{\mu}(x)\leq\delta_{0,x}+\frac{K_{C}}{L^{2}|\!|\!|x|\!|\!|_{L}^{d-2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}}\bigg{(}1\wedge\frac{(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{2}}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{4}(1-\mu)^{2}}\bigg{)}.$ (1.21)
Theorem 1.2 and Proposition 1.3 immediately give the same near-critical estimate for $G_{p}$.
Theorem 1.4.
Under the hypotheses of Theorem 1.2, if $\alpha\neq 2$ there is a constant $K>0$ independent of $L$ such that for all $p\in[\frac{1}{2},p_{c}]$ and $x\in\mathbb{Z}^{d}$,
$G_{p}(x)\leq\delta_{0,x}+\frac{K}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\bigg{(}1\wedge\frac{1}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{2(\alpha\wedge 2)}(p_{c}-p)^{2}}\bigg{)}.$ (1.22)
If $\alpha=2$, we instead have
$G_{p}(x)\leq\delta_{0,x}+\frac{K}{L^{2}|\!|\!|x|\!|\!|_{L}^{d-2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}}\bigg{(}1\wedge\frac{(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{2}}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{4}(p_{c}-p)^{2}}\bigg{)}.$ (1.23)
Proof of Theorem 1.4.
Since $G_{p}(0)=1$ for all our models, it suffices to prove the bound for $x\neq 0$. We consider $\alpha\neq 2$ first. Using Theorem 1.2 and Proposition 1.3 with $P$, for $x\neq 0$ we have
$G_{p}(x)=A_{p}S_{\mu_{p}}(x)\leq\frac{A_{p}K_{S}}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\bigg{(}1\wedge\frac{1}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{2(\alpha\wedge 2)}(1-{\mu_{p}})^{2}}\bigg{)}.$ (1.24)
From (1.19), we have $A_{p}=1+O(L^{-d})$ uniformly in $p$. Also, it follows from the uniform estimate on $P(x)/D(x)$ that $P$ obeys Assumption 1.1 with constants independent of $p$ (they can be chosen to be $1+O(L^{-(\ell-1)d})$ times the corresponding constants for $D$), and hence $K_{S}$ is also independent of $p$. We bound $A_{p}K_{S}\leq K$, and it remains to prove
$1-{\mu_{p}}\asymp p_{c}-p.$ (1.25)
For this, we first note the bound (1.24) at $p=p_{c}$ implies finiteness of the bubble diagram $B(p_{c})=G_{p_{c}}*G_{p_{c}}(0)$ for Ising and SAW and finiteness of the triangle diagram $T(p_{c})=G_{p_{c}}^{*3}(0)$ for percolation, since $d>d_{c}$. It follows that ${\chi(p)}\asymp(p_{c}-p)^{-1}$. Since summing equation (1.18) over $x\in\mathbb{Z}^{d}$ gives
${\chi(p)}=\sum_{x\in\mathbb{Z}^{d}}G_{p}(x)=A_{p}\sum_{x\in\mathbb{Z}^{d}}S_{\mu_{p}}(x)=\frac{A_{p}}{1-{\mu_{p}}}=\frac{1+O(L^{-d})}{1-{\mu_{p}}},$ (1.26)
we get (1.25), as desired.
The proof for $\alpha=2$ differs only at (1.24), where we would use (1.21) in Proposition 1.3 instead. The additional logarithmic term in the denominator of (1.21) makes the bubble or triangle diagram finite also at $d=d_{c}$, and we conclude the proof the same way. ∎
1.4 Lace expansion
The proof of Theorem 1.2 uses the lace expansion for the self-avoiding walk [6, 30], percolation [17], and the Ising model [28]. The lace expansion is a perturbative method that produces a convolution equation for $G_{p}$, and its convergence requires a small parameter, which is $L^{-1}$ in our spread-out setting and is $(d-d_{c})^{-1}$ for nearest-neighbour models. We refer to [30] for an introduction to the lace expansion method. For the self-avoiding walk, the lace expansion equation is
$G_{p}=\delta+pD*G_{p}+\Pi_{p}*G_{p}\qquad(0\leq p\leq p_{c}),$ (1.27)
where $\delta(x)=\delta_{0,x}$ is the Kronecker delta and $\Pi_{p}:\mathbb{Z}^{d}\to\mathbb{R}$ is an explicit but complicated function encoding the self-avoiding constraint. Indeed, if $\Pi_{p}$ is absent, the equation reduces to the renewal equation
$C_{p}=\delta+pD*C_{p}\qquad(0\leq p\leq 1)$ (1.28)
for the random walk two-point function $C_{p}$. For percolation and the Ising model, the lace expansion equation is different from (1.27), but it is a recent observation of [25] that the equation can be transformed into (1.27) in the spread-out setting (see Section 3.2).
Common to previous works [16, 8, 9, 25] based on the lace expansion, we employ a bootstrap argument in $x$ space to prove that
$\lvert\Pi_{p}(x)\rvert\lesssim\bigg{(}\frac{1}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\bigg{)}^{\ell}=\bigg{(}\frac{1}{L^{d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d-\alpha\wedge 2}}\bigg{)}^{\ell}\qquad(x\neq 0),$ (1.29)
where $\ell\in\{2,3\}$ is defined in (1.17), for $L$ sufficiently large. The bound is due to geometric considerations of the interaction. For example, for percolation, a main contribution to $\Pi_{p}(x)$ is from the event that 00 and $x$ are connected by two disjoint paths (doubly-connected), so the BK inequality implies that it decays at least twice ($\ell=2$) as fast as the two-point function $G_{p}(x)$. We do not have new things to say about how $\Pi_{p}(x)$ is bounded, but our bootstrap step is simpler than all previous works, since we do not need to use deconvolution theorems, thanks to the exact random walk representation (1.18).
1.5 Strategy of proof
We now present the strategy of the proof of Theorem 1.2 for $0<\alpha\leq 2$. The extension to all $\alpha\leq 2+(\ell-1)(d-d_{c})$ will be discussed near the end of Section 1.5.
The main idea is to define a new distribution $P$ proportional to $pD+\Pi_{p}$, then (1.27) becomes like (1.28), and we can conclude the theorem using the uniqueness of solutions to the convolution equation. If $D(x)$ decays too fast (e.g., nearest-neighbour), this strategy does not work because $\Pi_{p}(x)$ can take negative values. However, if $\alpha$ is small, $D(x)$ has a fat tail and might overtake any negative parts of $\Pi_{p}(x)$, to make $P(x)$ positive everywhere.111 We are inspired by a similar observation made on $(\delta+\Pi_{p})*D$ by Chen and Sakai in [9, Remark 3.7]. This is indeed the case when $\alpha\leq 2$. We combine the decay of $\Pi_{p}(x)$ in (1.29) with the assumed lower bound of $D(x)$ in (1.6), to get
$(pD+\Pi_{p})(x)\gtrsim\frac{1}{L^{d}}\bigg{(}\frac{1}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d+\alpha}}-\frac{O(L^{-(\ell-1)d})}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{\ell(d-\alpha)}}\bigg{)}\qquad(x\neq 0)$ (1.30)
uniformly in $p\in[\frac{1}{2},p_{c}]$. Since the exponents of $|\!|\!|\frac{x}{L}|\!|\!|_{1}$ satisfy
$\ell(d-\alpha)=d+\alpha+[(\ell-1)d-(\ell+1)\alpha]\geq d+\alpha$ (1.31)
when $d\geq\frac{\ell+1}{\ell-1}\alpha=d_{c}$ (recall (1.1) and (1.17)), we get
$(pD+\Pi_{p})(x)\gtrsim\frac{1}{L^{d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d+\alpha}}[1-O(L^{-(\ell-1)d})]\geq 0$ (1.32)
when $L$ is sufficiently large. This allows us to define a probability distribution $P$ by $P(0)=0$ and
$P(x)=\frac{pD(x)+\Pi_{p}(x)}{\lVert(pD+\Pi_{p})\mathds{1}_{x\neq 0}\rVert_{1}}=\frac{pD(x)+\Pi_{p}(x)}{p+\sum_{y\neq 0}\Pi_{p}(y)}\qquad(x\neq 0).$ (1.33)
Then, by decomposing $\Pi_{p}(x)=\Pi_{p}(0)\delta_{0,x}+\Pi_{p}(x)\mathds{1}_{x\neq 0}$ in equation (1.27), we get
$[1-\Pi_{p}(0)]G_{p}$ $=\delta+[(pD+\Pi_{p})\mathds{1}_{x\neq 0}]*G_{p}$
$=\delta+\underbrace{\bigg{(}\frac{p+\sum_{y\neq 0}\Pi_{p}(y)}{1-\Pi_{p}(0)}\bigg{)}}_{=:{\mu_{p}}}P*[1-\Pi_{p}(0)]G_{p}.$ (1.34)
Comparing to the random walk equation (1.28) with kernel $P$, once we prove that ${\mu_{p}}\in[0,1]$, uniqueness then yields
$[1-\Pi_{p}(0)]G_{p}=S_{\mu_{p}},$ (1.35)
where $S_{\mu}$ is the random walk two-point function generated by $P$. The desired (1.18) follows, with $A_{p}=[1-\Pi_{p}(0)]^{-1}$.
For $\alpha>2$, the strategy still works if $D(x)$ decays slower than $\Pi_{p}(x)$. By (1.29), this imposes
$\ell(d-2)\geq d+\alpha.$ (1.36)
Using the fact that $\ell(d_{c}-2)=d_{c}+2$, this is equivalent to the condition
$\alpha\leq\ell(d-2)-d-[\ell(d_{c}-2)]+(d_{c}+2)=2+(\ell-1)(d-d_{c}),$ (1.37)
which we assume.
The full proof of Theorem 1.2 is carried out in Section 3. As shown above, the key step is to prove (1.27) and (1.29) for all models. In the bootstrap argument there, we use an extended version of Proposition 1.3, which we prove first in Section 2.
2 Random walk two-point function
In this section, we prove a stronger version of Proposition 1.3 which also estimates random walks whose kernel $P(x)$ is close to $D(x)$. With $\mu\in[0,1]$, we write $S_{\mu}$ for the random walk two-point function generated by $P$, defined as in (1.9). By partitioning according to the length $\lvert\omega\rvert$, we have
$S_{\mu}(x)=\delta_{0,x}+\sum_{n=1}^{\infty}\mu^{n}P^{*n}(x),$ (2.1)
where $P^{*n}$ denotes the $n$-fold convolution of $P$ with itself. The following proposition implies Proposition 1.3, by taking $P=D$.
Proposition 2.1.
Let $\alpha>0$, $d>\alpha\wedge 2$, and let $D$ obey Assumption 1.1. There is a constant $K_{C}>0$ depending only on $K_{D}$ and $\Delta$ such that: Whenever $P:\mathbb{Z}^{d}\to[0,\infty)$ is a $\mathbb{Z}^{d}$-symmetric probability distribution such that $P(0)=0$,
$\frac{\frac{1}{2}K_{D}^{-1}L^{\alpha}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}\leq P(x)\leq\frac{2K_{D}L^{\alpha}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}\qquad(x\neq 0),$ (2.2)
and
$1-\hat{P}(k)\begin{cases}<2-\frac{1}{2}\Delta&(k\in(-\pi,\pi]^{d})\\ >\frac{1}{2}\Delta&(\lVert k\rVert_{\infty}>L^{-1}),\end{cases}$ (2.3)
we have
$S_{\mu}(x)\leq\delta_{0,x}+\frac{K_{C}}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\bigg{(}1\wedge\frac{1}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{2(\alpha\wedge 2)}(1-\mu)^{2}}\bigg{)}\qquad(\alpha\neq 2)$ (2.4)
or
$S_{\mu}(x)\leq\delta_{0,x}+\frac{K_{C}}{L^{2}|\!|\!|x|\!|\!|_{L}^{d-2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}}\bigg{(}1\wedge\frac{(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{2}}{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{4}(1-\mu)^{2}}\bigg{)}~{}\qquad(\alpha=2)$ (2.5)
for all $\mu\in[0,1]$ and $x\in\mathbb{Z}^{d}$.
We prove Proposition 2.1 by plugging the following estimates on $P^{*n}$ into (2.1).
Lemma 2.2.
Let $P$ be a $\mathbb{Z}^{d}$-symmetric probability distribution that obeys (2.2) and (2.3). There are constants depending only on $K_{D}$ and $\Delta$ such that for all $n\geq 1$ and all $x\in\mathbb{Z}^{d}$,
$\lVert P^{*n}\rVert_{\infty}$ $\lesssim\frac{1}{L^{d}}\times\begin{cases}n^{-d/(\alpha\wedge 2)}&(\alpha\neq 2)\\ (n\log\frac{\pi n}{2})^{-d/2}&(\alpha=2),\end{cases}$ (2.6)
$P^{*n}(x)$ $\lesssim\frac{nL^{\alpha\wedge 2}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha\wedge 2}}\times\begin{cases}1&(\alpha\neq 2)\\ \log|\!|\!|\frac{x}{L}|\!|\!|_{1}&(\alpha=2).\end{cases}$ (2.7)
Proof.
These bounds have been obtained for $P=D$ under Assumption 1.1 in [7, 8, 9]. The proof uses only the estimates (1.6)–(1.7) for $D$, so it applies equally to $P$ satisfying (2.2)–(2.3) and produces constants independent of $P$.
For (2.6), we first use the upper and lower bounds in (2.2) to get
$1-\hat{P}(k)\asymp(L\lvert k\rvert)^{\alpha\wedge 2}\times\begin{cases}1&(\alpha\neq 2)\\ \log\frac{\pi}{2L\lvert k\rvert}&(\alpha=2)\end{cases}$ (2.8)
for $\lvert k\rvert\leq L^{-1}$, from [7, (A.6)–(A.12)] (take $h(y)=|\!|\!|y|\!|\!|_{1}^{-(d+\alpha)}$ and $\ell=1$). Then, in conjunction with (2.3) and $\lVert P\rVert_{\infty}\lesssim L^{-d}$ from (2.2), from [7, (A.1)–(A.4)] (for $\alpha\neq 2$) or [9, Section 2.2] (for $\alpha=2$) we obtain (2.6). For (2.7), only the upper bound of $P(x)$ is needed, and the argument is given around [8, (1.21)]. This concludes the proof. ∎
Proof of Proposition 2.1.
We prove the bound for $S_{1}$ first. Since $S_{\mu}\leq S_{1}$, this will also give the “$1$” part of the minimum in (2.4)–(2.5) for $S_{\mu}$. We follow [8, (2.3)] (for $\alpha\neq 2$) and [9, (2.1)] (for $\alpha=2$). If $\alpha<2$, we set $N=|\!|\!|\frac{x}{L}|\!|\!|_{1}^{\alpha}=(|\!|\!|x|\!|\!|_{L}/L)^{\alpha}$. For $P^{*n}(x)$ in (2.1), we use (2.7) if $n\leq N$ and use (2.6) if $n\geq N$. This gives
$S_{1}(x)-\delta_{0,x}\lesssim\frac{L^{\alpha}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}\sum_{n=1}^{N}n+\frac{1}{L^{d}}\sum_{n=N}^{\infty}n^{-d/\alpha}\lesssim\frac{L^{\alpha}N^{2}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}+\frac{N^{1-d/\alpha}}{L^{d}}=\frac{2}{L^{\alpha}|\!|\!|x|\!|\!|_{L}^{d-\alpha}},$ (2.9)
as desired. The same proof works for $\alpha>2$ after replacing all $\alpha$ by $2$. If $\alpha=2$, we set $N=|\!|\!|\frac{x}{L}|\!|\!|_{1}^{2}/\log|\!|\!|\frac{x}{L}|\!|\!|_{1}$ instead, and we have
$S_{1}(x)-\delta_{0,x}$ $\lesssim\frac{L^{2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}}{|\!|\!|x|\!|\!|_{L}^{d+2}}\sum_{n=1}^{N}n+\frac{1}{L^{d}}\sum_{n=N}^{\infty}(n\log n)^{-d/2}$
$\lesssim\frac{L^{2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}N^{2}}{|\!|\!|x|\!|\!|_{L}^{d+2}}+\frac{N^{1-d/2}}{L^{d}(\log N)^{d/2}}\leq\frac{2}{L^{2}|\!|\!|x|\!|\!|_{L}^{d-2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}},$ (2.10)
as desired.
It remains to prove the other alternative in the minimum in (2.4)–(2.5) for $S_{\mu}$ when $\mu<1$. We again consider the case $\alpha<2$ first. Using (2.7) on all terms in (2.1), we have
$S_{\mu}(x)-\delta_{0,x}\lesssim\frac{L^{\alpha}}{|\!|\!|x|\!|\!|_{L}^{d+\alpha}}\sum_{n=1}^{\infty}n\mu^{n}=\frac{1}{L^{\alpha}|\!|\!|x|\!|\!|_{L}^{d-\alpha}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{2\alpha}}\frac{\mu}{(1-\mu)^{2}},$ (2.11)
and the desired bound follows from using $\mu\leq 1$ in the numerator. The same proof works for $\alpha>2$ after replacing all $\alpha$ by 2. If $\alpha=2$, we have an extra factor of $\log|\!|\!|\frac{x}{L}|\!|\!|_{1}$ from (2.7), so
$S_{\mu}(x)-\delta_{0,x}\lesssim\frac{L^{2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}}{|\!|\!|x|\!|\!|_{L}^{d+2}}\sum_{n=1}^{\infty}n\mu^{n}=\frac{1}{L^{2}|\!|\!|x|\!|\!|_{L}^{d-2}\log|\!|\!|\frac{x}{L}|\!|\!|_{1}}\frac{(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{2}\mu}{{|\!|\!|\frac{x}{L}|\!|\!|_{1}^{4}}(1-\mu)^{2}},$ (2.12)
and the desired result again follows from $\mu\leq 1$. This concludes the proof. ∎
3 Proof of Theorem 1.2
In this section, we prove Theorem 1.2. Throughout the section, we assume $d>d_{c}$ if $\alpha\neq 2$, and assume $d\geq d_{c}$ if $\alpha=2$. We also assume $D$ (for percolation or self-avoiding walk) or $J$ (for Ising) obeys Assumption 1.1. As discussed in Section 1.5, our main task is to prove the lace expansion equation (1.27) and the estimate (1.29) for $\Pi_{p}(x)$. We employ a bootstrap argument, similar to but simpler than the ones used in [16, 8, 9, 25].
3.1 Bootstrap argument
The bootstrap argument compares $G_{p}$ to the upper bound of $C_{1}$ given in Proposition 1.3. Let $K_{C}$ denote the constant of either (1.20) or (1.21), depending on $\alpha$. We define the bootstrap function $b:[\frac{1}{2},p_{c}]\to[0,\infty]$ by
$b(p)=p\vee\sup_{x\neq 0}\frac{G_{p}(x)}{K_{C}L^{-(\alpha\wedge 2)}|\!|\!|x|\!|\!|_{L}^{-(d-\alpha\wedge 2)}}\qquad(\alpha\neq 2)$ (3.1)
or by
$b(p)=p\vee\sup_{x\neq 0}\frac{G_{p}(x)}{K_{C}L^{-2}|\!|\!|x|\!|\!|_{L}^{-(d-2)}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{-1}}\qquad(\alpha=2).$ (3.2)
(Recall our convention that $|\!|\!|\frac{x}{L}|\!|\!|_{1}\geq\frac{\pi}{2}>1$.) The function $b(p)$ satisfies $b(1)\leq 1$ since $G_{1}\leq C_{1}$ [8, Lemma 2.3], and it is continuous in $p\in[\frac{1}{2},p_{c})$ by [8, (3.31)–(3.34)] (for $\alpha\neq 2$) or [9, (3.13)–(3.16)] (for $\alpha=2$). The proof of continuity uses a Simon–Lieb type inequality to show that $G_{p}(x)$ decays faster than the denominator of (3.1) or (3.2) when $p<p_{c}$. We will prove the implication
$b(p)\leq 3\quad\implies\quad b(p)\leq 2$ (3.3)
for all $p<p_{c}$, so that $b(p)$ cannot take any values in the interval $(2,3]$. Continuity then implies that $b(p)\leq 2$ for all $p<p_{c}$. Using the left-continuity of $G_{p}$ at $p_{c}$, we will get $b(p_{c})\leq 2$ also.
The a priori estimate $b(p)\leq 3$ on the bootstrap function gives a control on $G_{p}(x)$ and allows the lace expansion to be carried out. For the Ising model, from $p\leq b(p)\leq 3$ and (1.15) we also get $\beta\leq 3[1+O(L^{-2d})]\leq 4$, so that the $D$ defined in (1.14) obeys Assumption 1.1. We collect the lace expansion equation and the diagrammatic estimates in the following proposition, whose proof we postpone to Section 3.2.
Proposition 3.1 (Lace expansion).
There is an $L_{1}$ such that for all $L\geq L_{1}$, if $b(p)\leq 3$ then there exists a $\mathbb{Z}^{d}$-symmetric function $\Pi_{p}:\mathbb{Z}^{d}\to\mathbb{R}$ for which
$G_{p}=\delta+pD*G_{p}+\Pi_{p}*G_{p},$ (3.4)
and
$\lvert\Pi_{p}(x)\rvert\lesssim\frac{\delta_{0,x}}{L^{d}}+\bigg{(}\frac{1}{L^{d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d-\alpha\wedge 2}}\bigg{)}^{\ell}\times\begin{cases}1&(\alpha\neq 2)\\ (\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{-\ell}&(\alpha=2)\end{cases}$ (3.5)
with a constant independent of $p,L,x$.
We note that, by considering the two cases $0<\lvert x\rvert\leq L$ and $\lvert x\rvert>L$, (3.5) implies that
$\sum_{x\neq 0}\lvert\Pi_{p}(x)\rvert\lesssim\sum_{0<\lvert x\rvert\leq L}\frac{1}{L^{\ell d}}+\sum_{\lvert x\rvert>L}\Bigl{(}\frac{1}{L^{\alpha\wedge 2}\lvert x\rvert^{d-\alpha\wedge 2}}\Bigr{)}^{\ell}\lesssim\frac{1}{L^{(\ell-1)d}}.$ (3.6)
Given $b(p)\leq 3$, (3.4) and (3.5) allow us to carry out the strategy of Section 1.5 and define $P$.
Proposition 3.2.
Let $0<\alpha\leq 2+(\ell-1)(d-d_{c})$. There is an $L_{2}\geq L_{1}$ such that for all $L\geq L_{2}$, if $b(p)\leq 3$ then the distribution $P$ of (1.33) is well-defined and equation (1.5) holds. Moreover, we have
$A_{p},\,\frac{{\mu_{p}}}{p}=1+O(L^{-d}),\qquad\frac{P(x)}{D(x)},\,\frac{1-\hat{P}(k)}{1-\hat{D}(k)}=1+O(L^{-(\ell-1)d}),$ (3.7)
with constants uniform in $p\in[\frac{1}{2},p_{c}]$, in $x\neq 0$, and in $k\neq 0$.
Proof.
Let $L\geq L_{1}$ and $b(p)\leq 3$, so that we can apply Proposition 3.1. We have noted in (1.31) that
$\ell(d-\alpha)\geq d+\alpha\qquad(\alpha\leq 2,\ d\geq d_{c}),$ (3.8)
and in (1.37) that
$\ell(d-2)\geq d+\alpha\qquad(2<\alpha\leq 2+(\ell-1)(d-d_{c})),$ (3.9)
so we always have
$\ell(d-\alpha\wedge 2)\geq d+\alpha.$ (3.10)
Using (3.5) and the lower bound of $D(x)$ in (1.6), we have
$\lvert\Pi_{p}(x)\rvert\lesssim\frac{1}{L^{\ell d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{\ell(d-\alpha\wedge 2)}}\leq\frac{1}{L^{\ell d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d+\alpha}}\lesssim\frac{1}{L^{(\ell-1)d}}D(x)\qquad(x\neq 0).$ (3.11)
By restricting to $L$ large and $p\geq\frac{1}{2}$, we can therefore ensure $pD(x)+\Pi_{p}(x)\geq 0$ and define the probability distribution $P$ by (1.33). Equation (1.5) follows.
It remains to prove (3.7). For $A_{p}$ and ${\mu_{p}}/p$, we use the (3.5), (3.6), and $p\geq\frac{1}{2}$ to get
$A_{p}$ $=[1-\Pi_{p}(0)]^{-1}=1+O(L^{-d}),$ (3.12)
$\frac{{\mu_{p}}}{p}$ $=A_{p}\Big{(}1+\frac{1}{p}\sum_{y\neq 0}\Pi_{p}(y)\Big{)}=A_{p}[1+O(L^{-(\ell-1)d})]=1+O(L^{-d}).$ (3.13)
For $P(x)/D(x)$, we use the definition of $P$ in (1.33), $p\geq\frac{1}{2}$, and (3.6) to write
$P(x)=\frac{D(x)+\frac{1}{p}\Pi_{p}(x)}{1+\frac{1}{p}\sum_{y\neq 0}\Pi_{p}(y)}=[1+O(L^{-(\ell-1)d})][D(x)+\frac{1}{p}\Pi_{p}(x)]\qquad(x\neq 0).$ (3.14)
Then we divide the equation by $D(x)$ and get the desired result from (3.11). For the Fourier transform, we use $\mathbb{Z}^{d}$-symmetry and the uniform estimate on $P(x)/D(x)$ to get
$1-\hat{P}(k)=\sum_{x\neq 0}\Bigl{(}1+\frac{P(x)}{D(x)}-1\Bigr{)}D(x)(1-\cos(k\cdot x))=[1+O(L^{-(\ell-1)d})](1-\hat{D}(k)).$ (3.15)
This concludes the proof. ∎
Lemma 3.3 (Bootstrap).
There is an $L_{3}\geq L_{2}$ such that for all $L\geq L_{3}$, if $p\in[\frac{1}{2},p_{c})$ and $b(p)\leq 3$ then ${\mu_{p}}<1$ and $b(p)\leq 2$.
Proof.
Let $L\geq L_{2}$ and $p\in[\frac{1}{2},p_{c})$. If $b(p)\leq 3$, we can apply Proposition 3.1, sum equation (3.4) over $x\in\mathbb{Z}^{d}$, and then divide by ${\chi(p)}=\sum_{x\in\mathbb{Z}^{d}}G_{p}(x)<\infty$, to get
$1=\frac{1}{{\chi(p)}}+p+\sum_{x\in\mathbb{Z}^{d}}\Pi_{p}(x).$ (3.16)
Since ${\chi(p)}\in[1,\infty)$, we have $1>p+\sum_{x\in\mathbb{Z}^{d}}\Pi_{p}(x)$, which is equivalent to ${\mu_{p}}<1$ by definition (1.5). Using (3.5) and (3.6), from the same inequality we also get $p\leq 1+O(L^{-d})$.
To prove $b(p)\leq 2$, it remains to estimate $G_{p}$. We first use Proposition 3.2 to get equation (1.5), and then we observe that both $[1-\Pi_{p}(0)]G_{p}$ and the random walk two-point function $S_{\mu_{p}}$ generated by $P$ are $\ell^{1}$ solutions to the convolution equation
$S=\delta+{\mu_{p}}P*S.$ (3.17)
It then follows from the uniqueness of solutions (the Fourier transform of the difference of two solutions is 00) that $[1-\Pi_{p}(0)]G_{p}=S_{\mu_{p}}$ everywhere. We use Proposition 2.1 on $P$ to estimate $S_{\mu_{p}}$, and we verify the hypotheses (2.2)–(2.3) using (3.7), Assumption 1.1, and by restricting to $L$ sufficiently large. This gives
$[1-\Pi_{p}(0)]G_{p}(x)=S_{\mu_{p}}(x)\leq S_{1}(x)\leq\frac{K_{C}}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\times\begin{cases}1&(\alpha\neq 2)\\ (\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{-1}&(\alpha=2)\end{cases}$ (3.18)
uniformly in $x\neq 0$. Putting the above into the definition of $b(p)$, and using $\Pi_{p}(0)=O(L^{-d})$ , we obtain $b(p)\leq 1+O(L^{-d})\leq 2$ for $L$ sufficiently large. This concludes the proof. ∎
Proof of Theorem 1.2.
Set $L_{0}=L_{3}$. We recall the bootstrap function $b(p)$ is continuous in $p\in[\frac{1}{2},p_{c})$, satisfies $b(1)\leq 1$, and cannot take any values in $(2,3]$ when $p\in[\frac{1}{2},p_{c})$ by Lemma 3.3. Therefore, by continuity we must have $b(p)\leq 2$ for all $p\in[\frac{1}{2},p_{c})$. Proposition 3.2 and the uniqueness of solution argument given around (3.17) then give all desired results for $p\in[\frac{1}{2},p_{c})$.
For the remaining critical case, we first use the uniform bound $b(p)\leq 2$ established above and the left-continuity of $G_{p}(x)$ at $p=p_{c}$ [8, Lemma 2.2] to get
$G_{p_{c}}(x)=\lim_{p\to p_{c}^{-}}G_{p}(x)\leq\frac{2K_{C}}{L^{\alpha\wedge 2}|\!|\!|x|\!|\!|_{L}^{d-\alpha\wedge 2}}\times\begin{cases}1&(\alpha\neq 2)\\ (\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{-1}&(\alpha=2)\end{cases}$ (3.19)
for all $x\neq 0$. This implies $b(p_{c})\leq 2$, so Proposition 3.2 still applies and gives equation (1.5) at $p_{c}$, but we need to slightly modify the uniqueness argument and use the $L^{2}$ Fourier transform. We first observe that (3.19) implies $G_{p_{c}}\in\ell^{2}(\mathbb{Z}^{d})$, since $d>d_{c}$ (for $\alpha\neq 2$) or $d\geq d_{c}$ (for $\alpha=2$). Similarly, Proposition 2.1 implies $S_{1}\in\ell^{2}(\mathbb{Z}^{d})$ in these dimensions. We claim ${\mu_{p_{c}}}=1$. Assume this for a moment, by (1.5) we find that both $[1-\Pi_{p_{c}}(0)]G_{p_{c}}$ and $S_{1}$ are $\ell^{2}$ solutions to
$S=\delta+P*S,$ (3.20)
so we can use the uniqueness of $\ell^{2}$ solutions (the $L^{2}$ Fourier transform of the difference of two solutions is 00) to conclude $[1-\Pi_{p_{c}}(0)]G_{p_{c}}=S_{1}$, as desired.
It remains to prove ${\mu_{p_{c}}}=1$. We rearrange equation (3.4) at $p_{c}$ as
$(\delta-p_{c}D-\Pi_{p_{c}})*G_{p_{c}}=\delta.$ (3.21)
This allows us to apply Proposition A.1 with $F=\delta-p_{c}D-\Pi_{p_{c}}\in\ell^{1}(\mathbb{Z}^{d})$ and $G=G_{p_{c}}\in\ell^{2}(\mathbb{Z}^{d})$. Since $\sum_{x\in\mathbb{Z}^{d}}G_{p_{c}}(x)=\chi(p_{c})=\infty$, we obtain $\sum_{x\in\mathbb{Z}^{d}}F(x)=1-p_{c}-\sum_{x\in\mathbb{Z}^{d}}\Pi_{p_{c}}(x)=0$, which is equivalent to ${\mu_{p_{c}}}=1$ by definition (1.5). This concludes the proof. ∎
3.2 Proof of Proposition 3.1
It remains to prove Proposition 3.1. The desired equation (3.4) is the lace expansion equation for the self-avoiding walk [6, 30]. The bound (3.5) on $\Pi_{p}$ is known as a diagrammatic estimate, and it is proved by applying convolution estimates to lace expansion diagrams, first done in [16]. For the self-avoiding walk, Proposition 3.1 is proved in [8, Proposition 3.1] (for $\alpha\neq 2$) and [9, Lemma 3.6] (for $\alpha=2$) (both denote our $\Pi_{p}$ by $\pi_{p}$), by improving the convolution estimates of [16] using $|\!|\!|\cdot|\!|\!|_{L}$. We therefore focus on percolation and the Ising model.
The lace expansion equation for percolation [17] and the Ising model [28] takes the form
$G_{p}=h_{p}+pD*h_{p}*G_{p},$ (3.22)
where $h_{p}$ is a perturbation of the Kronecker delta $\delta(x)=\delta_{0,x}$. It is observed in [25] that equation (3.22) can be transformed into the desired (3.4) in the spread-out setting. The proof in [25] uses a Banach algebra (as in [5]), and the main idea is to convolve equation (3.22) with the convolution inverse of $h_{p}$, defined by the Neumann series
$h_{p}^{-1}=(\delta-(\delta-h_{p}))^{-1}=\sum_{n=0}^{\infty}(\delta-h_{p})^{*n}.$ (3.23)
This gives $h_{p}^{-1}*G_{p}=\delta+pD*G_{p}$, so
$G_{p}=\delta+pD*G_{p}+(\delta-h_{p}^{-1})*G_{p},$ (3.24)
which is (3.4) with $\Pi_{p}=\delta-h_{p}^{-1}$. We simply estimate the series (3.23), using the following lemma. Note that both $f_{1},f_{2}$ are integrable since $\theta>0$, so the upper bound on $f_{1}*f_{2}$ takes the same form.
Lemma 3.4.
Let $d,L\geq 1$, $\theta>0$, and $\phi,\psi\geq 0$. Suppose $f_{1},f_{2}:\mathbb{Z}^{d}\to\mathbb{R}$ obey the bound
$\lvert f_{1}(x)\rvert,\lvert f_{2}(x)\rvert\leq\delta_{0,x}+\frac{1}{L^{d+\phi}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d+\theta}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{\psi}}\qquad(x\in\mathbb{Z}^{d}).$ (3.25)
There there is a constant $C=C(d,\theta,\psi)>0$ such that
$\lvert(f_{1}*f_{2})(x)\rvert\leq C\bigg{(}\delta_{0,x}+\frac{1}{L^{d+\phi}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d+\theta}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{\psi}}\bigg{)}\qquad(x\in\mathbb{Z}^{d}).$ (3.26)
Proof.
For $j\in\{1,2\}$, define $g_{j}=f_{j}-f_{j}(0)\delta$, so that $g_{j}(0)=0$ and $g_{j}(x)=f_{j}(x)$ when $x\neq 0$. Since
$f_{1}*f_{2}=f_{1}(0)f_{2}(0)\delta+f_{1}(0)g_{2}+f_{2}(0)g_{1}+g_{1}*g_{2},$ (3.27)
and since $\lvert f_{j}(0)\rvert\leq 1+L^{-(d+\phi)}\leq 2$, we only need to estimate $g_{1}*g_{2}$. By the hypothesis, we have
$\lvert g_{1}*g_{2}(x)\rvert$ $\leq\sum_{y\in\mathbb{Z}^{d}}\frac{1}{L^{d+\phi}|\!|\!|\frac{x-y}{L}|\!|\!|_{1}^{d+\theta}(\log|\!|\!|\frac{x-y}{L}|\!|\!|_{1})^{\psi}}\frac{1}{L^{d+\phi}|\!|\!|\frac{y}{L}|\!|\!|_{1}^{d+\theta}(\log|\!|\!|\frac{y}{L}|\!|\!|_{1})^{\psi}}$
$=\frac{L^{2\theta}}{L^{2\phi}}\sum_{y\in\mathbb{Z}^{d}}\frac{1}{|\!|\!|x-y|\!|\!|_{L}^{d+\theta}(\log|\!|\!|\frac{x-y}{L}|\!|\!|_{1})^{\psi}}\frac{1}{|\!|\!|y|\!|\!|_{L}^{d+\theta}(\log|\!|\!|\frac{y}{L}|\!|\!|_{1})^{\psi}}$
$\leq\frac{L^{2\theta}}{L^{2\phi}}\frac{C_{d,\theta,\psi}L^{-\theta}}{|\!|\!|x|\!|\!|_{L}^{d+\theta}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{\psi}}=\frac{C_{d,\theta,\psi}}{L^{\phi}}\frac{1}{L^{d+\phi}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d+\theta}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{\psi}},$ (3.28)
where in the last line we used the convolution estimate [9, Lemma 3.5] with $a_{1}=b_{1}=d+\theta$ and $a_{2}=b_{2}=\psi$ (or use [8, Lemma 3.2(i)] if $\psi=0$). This gives the desired estimate, since $\phi\geq 0$. ∎
Proof of Proposition 3.1 for percolation and the Ising model.
It is proved in [8, Proposition 3.1] (for $\alpha\neq 2$) and [9, Lemma 3.6] (for $\alpha=2$) (both denote our $h_{p}$ by $\pi_{p}$) that, under $b(p)\leq 3$, equation (3.22) holds with a function $h_{p}$ that obeys
$\lvert h_{p}(x)-\delta_{0,x}\rvert\leq K_{h}\frac{\delta_{0,x}}{L^{d}}+K_{h}\bigg{(}\frac{1}{L^{d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{d-\alpha\wedge 2}}\bigg{)}^{\ell}\times\begin{cases}1&(\alpha\neq 2)\\ (\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{-\ell}&(\alpha=2),\end{cases}$ (3.29)
where $K_{h}>0$ is independent of $p,L,x$. For percolation, the proof is based on [16, Proposition 1.8]; for the Ising model, the proof is based on [29, Corollaries 5.5, 5.10, and 5.13].
Set $\phi=(\ell-2)d$, $\psi=\ell\mathds{1}_{\alpha=2}$, and
$\theta=\ell(d-\alpha\wedge 2)-d\geq\alpha\wedge 2>0,$ (3.30)
where the middle inequality is due to $d\geq d_{c}=\frac{\ell+1}{\ell-1}(\alpha\wedge 2)$. We first apply Lemma 3.4 to $f_{1}=f_{2}=\frac{L^{d}}{K_{h}}(h_{p}-\delta)$, then a simple induction gives
$\lvert(h_{p}-\delta)^{*n}(x)\rvert\leq C^{n-1}\bigg{(}\frac{K_{h}}{L^{d}}\bigg{)}^{n}\bigg{(}\delta_{0,x}+\frac{1}{L^{(\ell-1)d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{\ell(d-\alpha\wedge 2)}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{\psi}}\bigg{)}$ (3.31)
for all $n\geq 1$. By restricting to $L$ large enough such that $CK_{h}L^{-d}<1$, we can sum (3.31) over $n\geq 1$ to get
$\bigg{\lvert}\sum_{n=1}^{\infty}(\delta-h_{p})^{*n}(x)\bigg{\rvert}\leq\frac{K_{h}L^{-d}}{1-CK_{h}L^{-d}}\bigg{(}\delta_{0,x}+\frac{1}{L^{(\ell-1)d}|\!|\!|\frac{x}{L}|\!|\!|_{1}^{\ell(d-\alpha\wedge 2)}(\log|\!|\!|\frac{x}{L}|\!|\!|_{1})^{\psi}}\bigg{)}.$ (3.32)
It follows that the Neumann series (3.23) defining $h_{p}^{-1}$ converges, and from equation (3.24) we get the desired (3.4). The desired estimate on $\Pi_{p}=\delta-h_{p}^{-1}=-\sum_{n=1}^{\infty}(\delta-h_{p})^{*n}$ also follows from (3.32). This concludes the proof. ∎
Appendix A Characterisation of critical point
In this appendix, we develop a tool used in the proof of Theorem 1.2 to conclude that ${\mu_{p_{c}}}=1$, or equivalently, that
$1=p_{c}+\sum_{x\in\mathbb{Z}^{d}}\Pi_{p_{c}}(x).$ (A.1)
This characterisation of the critical point $p_{c}$ is traditionally obtained by taking the $p\to p_{c}^{-}$ limit of the subcritical version of (A.1) presented in (3.16). The traditional way, however, uses the left-continuity of $\Pi_{p}$ at $p_{c}$, which for some models is not so straightforward to prove (see [15, Appendix A] for percolation and self-avoiding walk). We work directly at $p_{c}$ and avoid using the left-continuity of $\Pi_{p}$.
We consider functions $F,G:\mathbb{Z}^{d}\to\mathbb{R}$ that satisfy
$(F*G)(x)=\delta_{0,x}\qquad(x\in\mathbb{Z}^{d}).$ (A.2)
If both $F,G\in\ell^{1}(\mathbb{Z}^{d})$ or if both are positive, by summing (A.2) over all $x\in\mathbb{Z}^{d}$ and using Fubini’s theorem, we have
$\sum_{x\in\mathbb{Z}^{d}}F(x)\sum_{x\in\mathbb{Z}^{d}}G(x)=1.$ (A.3)
The main result of this appendix is an extension of the above to the case where $G\in\ell^{2}\setminus\ell^{1}(\mathbb{Z}^{d})$.
Proposition A.1.
Let $d\geq 1$. Suppose $F\in\ell^{1}(\mathbb{Z}^{d})$, $G\in\ell^{2}(\mathbb{Z}^{d})$, $G\geq 0$, and suppose they satisfy equation (A.2). Then
$\sum_{x\in\mathbb{Z}^{d}}G(x)=\infty\quad\implies\quad\sum_{x\in\mathbb{Z}^{d}}F(x)=0.$ (A.4)
The proof uses the Bochner–Riesz operator in the Fourier space. We write $\mathbb{T}^{d}=(\mathbb{R}/2\pi\mathbb{Z})^{d}$ for the continuum torus, and we use the (inverse) Fourier transform
$\hat{f}(k)=\sum_{x\in\mathbb{Z}^{d}}f(x)e^{ik\cdot x}\quad(k\in\mathbb{T}^{d}),\qquad\check{u}(x)=\int_{\mathbb{T}^{d}}u(k)e^{-ik\cdot x}\frac{dk}{(2\pi)^{d}}\quad(x\in\mathbb{Z}^{d})$ (A.5)
for $f\in\ell^{1}(\mathbb{Z}^{d})$ and $u\in L^{1}({\mathbb{T}^{d}})$. For $R>0$, $\alpha\geq 0$, and $u\in L^{1}({\mathbb{T}^{d}})$, we define
$B_{R}^{\alpha}[u](k)=\sum_{\lvert x\rvert\leq R}\biggl{(}1-\frac{\lvert x\rvert^{2}}{R^{2}}\biggr{)}^{\alpha}\check{u}(x)e^{ik\cdot x}\qquad(k\in{\mathbb{T}^{d}}).$ (A.6)
The Bochner–Riesz operator $B_{R}^{\alpha}$ is an approximate identity:
Lemma A.2.
Let $d\geq 1$. If $\alpha>\frac{d-1}{2}$ and if the function $u\in L^{1}({\mathbb{T}^{d}})$ is continuous at $k_{0}\in{\mathbb{T}^{d}}$, then
$\lim_{R\to\infty}B_{R}^{\alpha}[u](k_{0})=u(k_{0}).$ (A.7)
Proof of Proposition A.1 assuming Lemma A.2.
We use the Fourier transform. The hypotheses imply that $\hat{F}$ is continuous on ${\mathbb{T}^{d}}$ and that $\hat{G}\in L^{2}({\mathbb{T}^{d}})\subset L^{1}({\mathbb{T}^{d}})$. From equation (A.2) we get $\hat{F}(k)\hat{G}(k)=1$ for almost every $k\in{\mathbb{T}^{d}}$. Our goal is to prove $\hat{F}(0)=0$. We prove by contradiction.
Suppose $\hat{F}(0)\neq 0$. Since $\hat{F}$ is continuous, the function $1/\hat{F}(k)$ is defined and continuous in a neighbourhood of 00. We apply Lemma A.2 with $u=1/\hat{F}=\hat{G}\in L^{1}({\mathbb{T}^{d}})$, $k_{0}=0$, and any choice of $\alpha>\frac{d-1}{2}$, to get
$\lim_{R\to\infty}B_{R}^{\alpha}[1/\hat{F}](0)=\frac{1}{\hat{F}(0)}.$ (A.8)
On the other hand, since the inverse Fourier transform of $1/\hat{F}=\hat{G}$ is $G$, the definition of $B_{R}^{\alpha}$ gives
$B_{R}^{\alpha}[1/\hat{F}](0)=B_{R}^{\alpha}[\hat{G}](0)=\sum_{\lvert x\rvert\leq R}\biggl{(}1-\frac{\lvert x\rvert^{2}}{R^{2}}\biggr{)}^{\alpha}G(x)\to\sum_{x\in\mathbb{Z}^{d}}G(x)=\infty,$ (A.9)
by monotone convergence. This is a contradiction to $\hat{F}(0)\neq 0$, so the proof is complete. ∎
It remains to prove Lemma A.2. The lemma is a slight variant of [14, Proposition 4.1.9], which states that $B_{R}^{\alpha}[u]$ converges uniformly to $u$ if $u$ is continuous on all of ${\mathbb{T}^{d}}$. In that proof, it is shown that, when $\alpha>\frac{d-1}{2}$, there is a family of kernels $L^{\alpha,R}:{\mathbb{T}^{d}}\to\mathbb{C}$ for which
$B_{R}^{\alpha}[u]=u*L^{\alpha,R}\qquad(u\in L^{1}({\mathbb{T}^{d}})),$ (A.10)
with the properties that222 Properties (A.11)–(A.13) make the family $\{L^{\alpha,R}\}_{R>0}$ an approximate identity, also known as a summability kernel.
[…]
for all $R>0$ and all $0<\delta<\pi$. ([14] uses the torus with period $1$ rather than our $2\pi$.) The proof of (A.13) also yields the stronger statement that
$\sup_{\delta\leq\lVert k\rVert_{\infty}\leq\pi}\lvert L^{\alpha,R}(k)\rvert\leq C_{d,\alpha,\delta}^{\prime}R^{-(\alpha-\frac{d-1}{2})}$ (A.14)
for all $R>0$.
Proof of Lemma A.2.
We modify the proof of [14, Theorem 1.2.19]. Let $u\in L^{1}({\mathbb{T}^{d}})$ be continuous at $k_{0}\in{\mathbb{T}^{d}}$, so that for any $\varepsilon>0$ there is a $\delta>0$ for which $\lVert k\rVert_{\infty}<\delta$ implies $\lvert u(k_{0}-k)-u(k_{0})\rvert<\varepsilon$. We use (A.10) and (A.12) to write
[…]
By splitting the integral according to how $\lVert k\rVert_{\infty}$ compares with $\delta$, and by using (A.14), we get
$\lvert B_{R}^{\alpha}[u](k_{0})-u(k_{0})\rvert\leq\varepsilon\lVert L^{a,R}\rVert_{1}+O(R^{-(\alpha-\frac{d-1}{2})})\big{(}\lVert u\rVert_{1}+(2\pi)^{d}\lvert u(k_{0})\rvert\big{)}.$ (A.16)
Sending $R\to\infty$, (A.11) then gives
$\limsup_{R\to\infty}\,\lvert B_{R}^{\alpha}[u](k_{0})-u(k_{0})\rvert\leq C_{d,\alpha}\varepsilon,$ (A.17)
and we are done because $\varepsilon>0$ is arbitrary. ∎
Acknowledgements
The work was supported in part by NSERC of Canada. We thank Noe Kawamoto, Romain Panis, and Gordon Slade for comments on a preliminary version.